OpenStudy (anonymous):

Find the slope of each of the three line segments in the graph below. Please show all of your work. Between which two points is the slope the smallest? points are on 3,3 7,6 9,9

OpenStudy (anonymous):

m1=(6-3)/(7-3) = 3/4 m2=(9-3)/(9-3) = 1 m3=(9-6)/(9-7)=3/2 so, the smallest slope is m1 which is between points: (3,3) & (7,6)

OpenStudy (anonymous):

inik that make sense ill go with that! thank u